# Reduction of Cosymplectic groupoids by cosymplectic moment maps

### v2

@@ Abstract.
Acknowledgments: We are grateful to David Iglesias for his valuable help in the discussions that inspired this work. We thank Cristian Ortiz for helpful and insightful discussions on the reduction of Poisson groupoids. We would like to thank Javier de Lucas and Xavier Rivas for inviting us to their Gamma seminar, where many ideas from this article were improved. López Garcia was supported by grant # 2022/04705-8, São Paulo Research Foundation (FAPESP).
@@ 1. Introduction
Cosymplectic manifolds arise as the odd counterpart of symplectic manifolds [22]. They can also be found as the underlying structures of other types of geometry as CR geometry [14], the odd-dimensional counterpart of Kähler manifolds and cokähler structures [2], almost contact manifolds [3], and in the description of time-dependent mechanics [19] and classical field theories [11]. Recently, these structures have been studied in the context of Lie groupoids giving rise to cosymplectic groupoids and their respective infinitesimal counterpart [13, 15] as central extensions of the underlying Poisson structure.
In the literature, several results about symmetries and reductions for geometric structures exist. In the case of cosymplectic geometry, we can consider two of them: the reduction of the associated Poisson structure and the Albert reduction using cosymplectic moment maps (analogous to the Marsden-Weinstein-Meyer symplectic reduction). The key difference between them is that the Poisson reduction does not preserve the nondegenerate condition in the cosymplectic case. On the other hand, the quotient of the level set of the cosymplectic moment map becomes cosymplectic again. For symplectic groupoids and Poisson structures on the units, these symmetries and reductions can be related by the integration theory of Lie groupoids [17]. The study of symmetries and reduction was later extended to Lie algebroids equipped with infinitesimal multiplicative 2-forms [6]. It is important to note that in [8], the case of IM 2-form reductions for quotients of vector bundles has been studied; in particular, Poisson structure quotients are studied. In contrast, our results focus on moment map reductions targeting co-symplectic structure reductions.
These notes extend the reduction technique in [6] to the particular case of cosymplectic groupoids and their Lie algebroids, taking into account the results in [15] and adapting the cosymplectic reduction theorem as in [1]. In Section 2 we start with an overview of cosymplectic manifolds, groupoids, and their infinitesimal data. In Section 3 we present the first main result of this article, which is a cosymplectic reduction by using a cosymplectic moment map in the context of Lie groupoids. Indeed, we consider symmetries of the cosymplectic forms equipped with a cosymplectic moment map that is also a groupoid morphism. Under certain conditions, the quotient of the 0-level set by the group action turns out to be a cosymplectic groupoid again (see Theorem 3.4).
@@
Section 4.2 investigates how the cosymplectic reduction of a cosymplectic groupoid is related to the symplectic reduction of a particular symplectic subgroupoid. On a cosymplectic groupoid, there is a symplectic foliation on the space of arrows. The leaf containing the units is a symplectic subgroupoid. In Theorem 4.8, we show how cosymplectic reduction induces a Marsden-Weinstein symplectic reduction on the previously mentioned leaf. In section 4.3, using the relation between cosymplectic groupoids and $S^{1}$-central extensions with vanishing multiplicative Chern class, we study what happens to the multiplicative Chern class of the reduced groupoid. In Theorem 4.11, we consider an oversymplectic groupoid with an action of a compact group; if the multiplicative Chern class of the initial groupoid vanishes, then the same is true for the reduced groupoid.
@@ 2. Preliminaries
This section is dedicated to presenting the main objects of this article. Since we study the symmetries of cosymplectic groupoids and their infinitesimal counterpart, we present all the necessary definitions and tools to be used here..
2.1. Fundamental concepts about cosymplectic structures.
To establish the notation and as a quick reference we describe the relation between a cosymplectic manifold and its Poisson structure; for a more detailed description the reader can consult [15].
@@
Let $(M,\omega)$ be a symplectic manifold of dimension $2m$. Let $i:N\hookrightarrow M$ be submanifold and $X$ be a symplectic vector field transverse to $N$. Then, $(N,i^{*}\omega,i^{*}(\imath_{X}\omega))$ is a cosymplectic manifold, in fact $d(i^{*}\omega)=0$, $di^{*}(\imath_{X}\omega))=i^{*}(\mathcal{L}_{X}\omega-\imath_{X}d\omega)=0$, and because $X$ is traverse to $N$, $\dim(N)=2(m-1)+1$ and $(i^{*}\omega)^{m-1}\wedge i^{*}(\imath_{X}\omega))$ is a volume form. Thus, the Lichnerowicz isomorphism is given by $\flat(Y)=i^{*}(\imath_{\tilde{Y}}\omega+(\imath_{\tilde{Y}}\imath_{X}\omega)\imath_{X}\omega),$ where $Y$ is $i$-related to $\tilde{Y}\in\mathfrak{X}(M)$.
Remark 2.2.
The converse is also true and was originally stated in [1, Prop 2]. The statement is as follows: if a manifold $Q$ is equipped with closed forms $\omega\in\Omega^{2}(Q)$ and $\eta\in\Omega^{1}(M)$ for which $\flat$ in (1) is an isomorphism and there exists a vector field as in (2), then $Q$ is odd-dimensional and $(Q,\omega,\eta)$ is a cosymplectic manifold. $\diamond$
@@
Furthermore, the Poisson structure of Example 2.3 is a particular case of a more general result. It is known that $\ker(\eta)\subset TQ$ defines a regular integrable foliation $\mathcal{F}$ of corank 1 [15]. Moreover, the leaves of $\mathcal{F}$ are symplectic with respect to the restriction of $\omega$. Indeed, there is a one to one correspondence between cosymplectic structures on $Q$ and regular Poisson structures of corank 1 on $Q$ with a Poisson vector field transverse to the symplectic foliation (see [18]).
@@
2.2. Symmetries and reduction
For this section, we assume a Lie group $G$ acting on a cosymplectic manifold on $(Q,\omega,\eta)$ by cosymplectic diffeomorphisms. Thus, it induces a $G$ Poisson action. In the case of a free and proper action, we obtain a reduced Poisson structure $\pi_{r}$ on $Q/G$ with quotient map $q:Q\to Q/G$ as Poisson morphism, that is $q_{*}\pi_{Q}=\pi_{r}$. By the correspondence between regular Poisson structures of corank 1 endowed with a Poisson vector field transverse to the symplectic foliation and cosymplectic structures, we have that a Poisson $Q/G$ reduction can be non-cosymplectic due to issues of rank or transversality. Nevertheless, there is a reduction procedure that preserve the cosymplectic structure, which is called cosymplectic reduction (see [1, 9]).
Let $G$ be a Lie group with Lie algebra $\mathfrak{g}=Lie(G)$. Let $(Q,\omega,\eta)$ be a cosymplectic manifold with an action of $G$ by cosymplectic morphisms. A smooth function $J:Q\to\mathfrak{g}^{*}$ is called a cosymplectic moment map if it is $G$-equivariant with respect to the coadjoint action, and for any $v\in\mathfrak{g}$ the Hamiltonian vector field $X_{\langle J,v\rangle}$ is the infinitesimal generator $v_{Q}\in\mathfrak{X}(Q)$. The value $0\in\mathfrak{g}^{*}$ is called a clean value if $J^{-1}(0)\subset Q$ is a submanifold and $\ker(dJ)_{x}=T_{x}J^{-1}(0)$ for all $x\in J^{-1}(0)$ (in [1, 23], it is called a weakly regular value). Clearly, a regular value is a clean value. Under these conditions there is a cosymplectic version of the classical Marsden-Weinstein reduction theorem [23]:
Theorem 2.4 ([1]).
Let $G$ be a Lie group acting on a cosymplectic manifold $(Q,\omega,\eta)$ by cosymplectic automorphisms. Let $J:Q\to\mathfrak{g}^{*}$ be a cosymplectic moment map such that $\xi(\langle J,v\rangle)=0$ for all $v\in\mathfrak{g}$. If 00 is a clean value of $J$, and $G$ acts freely and properly on $J^{-1}(0)$, then $Q_{red}=J^{-1}(0)/G$ is a cosymplectic manifold. The cosymplectic structure $(\omega_{red},\eta_{red})$ is defined by
@@ 2.2. Symmetries and reduction
Similarly to the symplectic case, when the action of $G$ on $Q$ is free, it follows that 0 is a regular value of $J$. From the Lichnerowicz isomorphism we observe that $v_{Q}=\flat^{-1}(d\langle J,v\rangle)$. Thus, $d\langle J_{x},v\rangle\neq 0$ if and only if $v_{Q}(x)\neq 0$ along $J^{-1}(0)$. Since $v_{Q}(e^{\tau v}x)=\left.\frac{d}{dt}\right|_{t=\tau}e^{tv}x$, then by the freeness of the action we conclude that $v_{Q}(x)\neq 0$ for any $v\neq 0\in\mathfrak{g}$. Consequently $dJ_{x}:T_{x}Q\to\mathfrak{g}^{*}$ is a surjective map. $\diamond$
Using a correspondence between cosymplectic moment maps on $Q$ and symplectic moment maps on $Q\times R$, we can study conditions under which cosymplectic moment maps exist.
@@
Let $i:Q\to Q\times\mathbb{R}$ the inclusion, and $p_{Q}$ and $p_{\mathbb{R}}$ be the canonical projections from $Q\times\mathbb{R}$ to $Q$ and $\mathbb{R}$, respectively. Consider $Q\times\mathbb{R}$ with the symplectic form $\tilde{\omega}=p_{Q}^{*}\omega+p_{\mathbb{R}}^{*}dt\wedge p_{Q}^{*}\eta$. There is a trivial extension of the $G$-action on $Q$ to a $G$-action on $Q\times\mathbb{R}$ given by $g\cdot(q,r)=(g\cdot q,r)$. We claim that there exists a one-to-one correspondence between cosymplectic moment maps on $(Q,\omega,\eta)$ satisfying $\xi\langle J,v\rangle=0$, and symplectic moment maps on $(Q\times\mathbb{R},\tilde{\omega})$.
Note that the vector fields $v_{Q}$ and $v_{Q\times\mathbb{R}}$ are $p_{Q}$-related. Thus, if $J:Q\to\mathfrak{g}^{*}$ is a cosymplectic moment map such that $\xi\langle J,v\rangle=0$, then $\tilde{J}:\mathbb{Q}\times\mathbb{R}\to\mathfrak{g}^{*}$ given by $\tilde{J}(q,r)=J(q)$ satisfies
@@
On the other hand, given a symplectic moment map $\tilde{J}:Q\times\mathbb{R}\to\mathfrak{g}^{*}$, we verify $\partial_{r}\langle\tilde{J},v\rangle=\imath_{\partial_{r}}\imath_{v_{Q\times\mathbb{R}}}\tilde{\omega}=p_{Q}^{*}(\imath_{v_{Q}}\eta)=0.$ Thus, $\tilde{J}$ does not depend on $r$, and $J:=i^{*}\tilde{J}$ is well-defined. If $\tilde{\xi}=i_{*}(\xi)$, then $\imath_{\tilde{\xi}}\imath_{v_{Q\times\mathbb{R}}}\tilde{\omega}=0.$ Therefore, $\xi\langle J,v\rangle=\tilde{\xi}\langle\tilde{J},v\rangle=0.$ Consequently, from this correspondence we conclude that the obstruction to the existence of $\tilde{J}$ is the same as $J$. The result follows from [24, Prop. 4.5.17] and from the homotopy equivalence between $Q$ and $Q\times\mathbb{R}$. ∎
2.3. Cosymplectic groupoids and their infinitesimal counterpart.
The concept of cosymplectic groupoid was originally introduced in [13], and have recently been studied in [15]. This structure is defined as a Lie groupoid $\mathcal{G}\rightrightarrows M$ endowed with multiplicative differential forms $\omega\in\Omega_{cl}^{2}(\mathcal{G})$ and $\eta\in\Omega_{cl}^{1}(\mathcal{G})$, such that $(\mathcal{G},\omega,\eta)$ is a cosymplectic manifold.
Since the differential forms are multiplicative, the Poisson structure $(\mathcal{G},\pi_{\mathcal{G}})$ is multiplicative and the Reeb vector field $\xi$ is bi-invariant. In addition, it is possible to prove that the dimension of $\mathcal{G}$ is $2$dim$(M)+1$. On the other hand, let $\Sigma$ be the leaf of the symplectic foliation on $\mathcal{G}$ containing $M\hookrightarrow\mathcal{G}$. The groupoid $(\Sigma,\omega|_{\Sigma})\rightrightarrows M$ is symplectic and it integrates the Poisson structure $(M,\pi_{M})$, where $\pi_{M}=t_{*}\pi_{\mathcal{G}}$. For a detailed discussion, the reader may refer to [15].
Following the two canonical constructions of cosymplectic manifolds in Examples 2.1 and 2.3, as a corank 1 submanifold and the product by a 1-dimensional manifold, it is possible to produce examples of cosymplectic groupoids.
Example 2.7.
Let $(\mathcal{G},\omega)\rightrightarrows M$ be a symplectic groupoid. Consider $i:\mathcal{G}_{0}\hookrightarrow\mathcal{G}$ a Lie subgroupoid over $M_{0}\hookrightarrow M$, both manifolds of codimension 1. If there exists a multiplicative symplectic vector field $X$ transverse to $\mathcal{G}_{0}$, then $(\mathcal{G}_{0},i^{*}\omega,i^{*}(\imath_{X}\omega))\rightrightarrows M$ is a cosymplectic groupoid. For a concrete example, consider a manifold $Q$, $\mathcal{G}=T^{*}(Q\times\mathbb{R})$ and $\omega=\omega_{can}$. Let $\mathcal{G}_{0}$ be the product of $T^{*}Q\rightrightarrows Q$ and ${}\mathbb{R}\rightrightarrows\{*\}$. If $r$ is the coordinate of the base of $T^{*}\mathbb{R}\to\mathbb{R}$, then the vector field $X=\partial r$ is multiplicative, symplectic and transverse to $\mathcal{G}_{0}$.
Example 2.8.
Associated with the construction of Example 2.3 we can consider two product groupoids. Let $(\mathcal{G}_{0},\omega)\rightrightarrows M$ be a symplectic groupoid. Consider $\mathcal{G}_{1}=\mathcal{G}_{0}\times S^{1}\rightrightarrows M$ and $\mathcal{G}_{2}=\mathcal{G}_{0}\times S^{1}\rightrightarrows M\times S^{1}$, where the structural maps of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ coming from the groupoids $S^{1}\rightrightarrows\{*\}$ and $S^{1}\rightrightarrows S^{1}$, respectively. The cosymplectic forms $pr_{\mathcal{G}_{0}}^{*}\omega$ and $pr_{S^{1}}d\theta$ are multiplicative in $\mathcal{G}_{1}$ making it a cosymplectic groupoid. However, $(\mathcal{G}_{2},\text{pr}_{\mathcal{G}_{0}}^{*}\omega,\text{pr}_{S^{1}}^{*}d\theta)$ is not a cosymplectic groupoids by dimensional issues.
We now proceed to describe the infinitesimal counterpart of these groupoids. For this, we recall the theory of infinitesimal multiplicative forms, as in [5]. These structures are defined on the Lie algebroid $A=Lie(\mathcal{G})$. For any multiplicative form $\alpha\in\Omega_{cl}^{k}(\mathcal{G})$ we define the vector-bundle map $\mu:A\to\Lambda^{k-1}T^{*}M$ covering the identity, defined by
where $u\in A$ and $X_{i}\in TM$. This definition can also be written as $\epsilon^{*}(\imath_{u^{R}}\alpha)=\mu(u)$, where $u^{R}$ is the right invariant vector field on $\mathcal{G}$ induced by $u\in A$. The bundle map $\mu$ is called IM k-form and it satisfies the compatibility conditions

### v3

@@ Abstract.
Acknowledgments: We are grateful to David Iglesias for his valuable help in the discussions that inspired this work. We thank Cristian Ortiz for helpful and insightful discussions on the reduction of Poisson groupoids. We would like to thank Javier de Lucas and Xavier Rivas for inviting us to their Gamma seminar, where many ideas from this article were improved. We are grateful to Luca Vitagliano for his detailed explanations on cosymplectic structures via line bundles and their associated cosymplectic groupoids. We thank the anonymous referee for valuable comments and detailed suggestions that significantly improved the presentation of this article. López Garcia was partially supported by grant # 2022/04705-8, São Paulo Research Foundation (FAPESP).
@@ 1. Introduction
Cosymplectic manifolds arise as an odd counterpart of symplectic manifolds [Li]. They can also be found as the underlying structures of other types of geometry as CR geometry [DT], the odd-dimensional counterpart of Kähler manifolds and co-Kähler structures [Bl1], almost contact manifolds [Bl0], and in the description of time-dependent mechanics [JLucas, GM] and classical field theories [LSM]. Recently, these structures have been studied in the context of Lie groupoids giving rise to cosymplectic groupoids and their respective infinitesimal counterpart [AW, FI] as central extensions of the underlying Poisson structure.
In the literature, several results about symmetries and reductions for geometric structures exist. In the case of cosymplectic geometry, we can consider two of them: the reduction of the associated Poisson structure and the Albert reduction using cosymplectic moment maps (analogous to the Marsden-Weinstein-Meyer symplectic reduction). The key difference between them is that the Poisson reduction does not preserve the nondegenerate condition in the cosymplectic case. On the other hand, the quotient of the level set of the cosymplectic moment map becomes cosymplectic again. For symplectic groupoids and Poisson structures on the units, these symmetries and reductions can be related by the integration theory of Lie groupoids [FOR]. The study of symmetries and reduction was later extended to Lie algebroids equipped with infinitesimal multiplicative 2-forms [BC, BC2]. It is important to note that in [CO], the case of IM 2-form reductions for quotients of vector bundles has been studied; in particular, Poisson structure quotients are studied. In contrast, our results focus on moment map reductions targeting co-symplectic structure reductions.
These notes extend the reduction technique in [BC2] to the particular case of cosymplectic groupoids and their Lie algebroids, taking into account the results in [FI] and adapting the cosymplectic reduction theorem as in [Al, JLucas]. In Section 2 we start with an overview of cosymplectic manifolds, groupoids, and their infinitesimal data. In Section 3 we present the first main result of this article, which is a cosymplectic reduction by using a cosymplectic moment map in the context of Lie groupoids. Indeed, we consider symmetries of the cosymplectic forms equipped with a cosymplectic moment map that is also a groupoid morphism. Under certain conditions, the quotient of the 0-level set by the group action turns out to be a cosymplectic groupoid again (see Theorem 3.4).
@@
Section 4.2 investigates how the cosymplectic reduction of a cosymplectic groupoid is related to the symplectic reduction of a particular symplectic subgroupoid. On a cosymplectic groupoid, there is a symplectic foliation on the space of arrows. The leaf containing the units is a symplectic subgroupoid. In Theorem 4.8, we show how cosymplectic reduction induces a Marsden-Weinstein symplectic reduction on the previously mentioned leaf. In section 4.3, using the relation between cosymplectic groupoids and $S^{1}$-central extensions with vanishing multiplicative Chern class, we study what happens to the multiplicative Chern class of the reduced groupoid. In Theorem 4.11, we consider an oversymplectic groupoid with an action of a compact, connected Lie group; if the multiplicative Chern class of the initial groupoid vanishes, then the same is true for the reduced groupoid.
@@ 2. Preliminaries
This section is dedicated to presenting the main objects of this article. Since we study the symmetries of cosymplectic groupoids and their infinitesimal counterpart, we present all the necessary definitions and tools to be used here.
2.1. Cosymplectic structures
To establish the notation and as a quick reference, we describe the relation between a cosymplectic manifold and its Poisson structure; for a more detailed description, the reader can consult [FI].
@@ 2.1. Cosymplectic structures
Remark 2.1.
The converse is also true and was originally stated in [Al, Prop 2]. The statement is as follows: if a manifold $Q$ is equipped with closed forms $\omega\in\Omega^{2}(Q)$ and $\eta\in\Omega^{1}(M)$ for which $\flat$ in (1) is an isomorphism and there exists a vector field as in (2), then $Q$ is odd-dimensional and $(Q,\omega,\eta)$ is a cosymplectic manifold. $\diamond$
@@
Let $(M,\omega)$ be a symplectic manifold of dimension $2m$. Let $i:N\hookrightarrow M$ be submanifold and $X$ be a symplectic vector field transverse to $N$. Then, $(N,i^{*}\omega,i^{*}(\imath_{X}\omega))$ is a cosymplectic manifold, in fact $d(i^{*}\omega)=0$, $di^{*}(\imath_{X}\omega)=i^{*}(\mathcal{L}_{X}\omega-\imath_{X}d\omega)=0$, and because $X$ is traverse to $N$, $\dim(N)=2(m-1)+1$ and $(i^{*}\omega)^{m-1}\wedge i^{*}(\imath_{X}\omega))$ is a volume form. The Lichnerowicz isomorphism is given by $\flat(Y)=i^{*}(\imath_{\tilde{Y}}\omega+(\imath_{\tilde{Y}}\imath_{X}\omega)\imath_{X}\omega),$ where $Y$ is $i$-related to $\tilde{Y}\in\mathfrak{X}(M)$.
@@
Furthermore, the Poisson structure of Example 2.3 is a particular case of a more general result. It is clear that $\ker(\eta)\subset TQ$ defines a regular integrable foliation $\mathcal{F}$ of corank 1. Moreover, the leaves of $\mathcal{F}$ are symplectic with respect to the restriction of $\omega$. Indeed, there is a one to one correspondence between cosymplectic structures on $Q$ and regular Poisson structures of corank 1 on $Q$ with a Poisson vector field transverse to the symplectic foliation (see [GMP]).
@@
2.1.1. Symmetries and reduction
For this section, we assume a Lie group $G$ acting on a cosymplectic manifold on $(Q,\omega,\eta)$ by cosymplectic diffeomorphisms. Thus, it induces a $G$ Poisson action. In the case of a free and proper action, we obtain a reduced Poisson structure $\pi_{r}$ on $Q/G$ with quotient map $q:Q\to Q/G$ as Poisson morphism, that is $q_{*}\pi_{Q}=\pi_{r}$. By the correspondence between regular Poisson structures of corank 1 endowed with a Poisson vector field transverse to the symplectic foliation and cosymplectic structures, we have that a Poisson $Q/G$ reduction can be non-cosymplectic due to issues of rank or transversality. Nevertheless, there is a reduction procedure that preserves the cosymplectic structure, which is called cosymplectic reduction (see [Al, CNY]).
Let $G$ be a Lie group with Lie algebra $\mathfrak{g}=Lie(G)$. Let $(Q,\omega,\eta)$ be a cosymplectic manifold with an action of $G$ by cosymplectic morphisms. A smooth function $J:Q\to\mathfrak{g}^{*}$ is called a cosymplectic moment map if it is $G$-equivariant with respect to the coadjoint action, and for any $v\in\mathfrak{g}$ the Hamiltonian vector field $X_{\langle J,v\rangle}$ is the infinitesimal generator $v_{Q}\in\mathfrak{X}(Q)$. The value $0\in\mathfrak{g}^{*}$ is called a clean value if $J^{-1}(0)\subset Q$ is a submanifold and $\ker(dJ)_{x}=T_{x}J^{-1}(0)$ for all $x\in J^{-1}(0)$ (in [Al, MW], it is called a weakly regular value). Clearly, a regular value is a clean value. Under these conditions there is a cosymplectic version of the classical Marsden-Weinstein reduction theorem [MW]:
Theorem 2.4 ([Al]).
Let $G$ be a Lie group acting on a cosymplectic manifold $(Q,\omega,\eta)$ by cosymplectic automorphisms. Let $J:Q\to\mathfrak{g}^{*}$ be a cosymplectic moment map such that $\xi(\langle J,v\rangle)=0$ for all $v\in\mathfrak{g}$. If 0 is a clean value of $J$, and $G$ acts freely and properly on $J^{-1}(0)$, then $Q_{red}=J^{-1}(0)/G$ is a cosymplectic manifold. The cosymplectic structure $(\omega_{red},\eta_{red})$ is defined by
@@ 2.1.1. Symmetries and reduction
Similarly to the symplectic case, when the action of $G$ on $Q$ is free, it follows that 0 is a regular value of $J$. From the Lichnerowicz isomorphism we observe that $v_{Q}=\flat^{-1}(d\langle J,v\rangle)$. By the freeness of the action we conclude that $v_{Q}(x)\neq 0$ for any $v\neq 0\in\mathfrak{g}$. Consequently $dJ_{x}:T_{x}Q\to\mathfrak{g}^{*}$ is a surjective map. $\diamond$
Using a correspondence between cosymplectic moment maps on $Q$ and symplectic moment maps on $Q\times\mathbb{R}$, we can study conditions under which cosymplectic moment maps exist.
@@
[…]
Note that the vector fields $v_{Q}$ and $v_{Q\times\mathbb{R}}$ are $p_{Q}$-related. Thus, if $J:Q\to\mathfrak{g}^{*}$ is a cosymplectic moment map such that $\xi\langle J,v\rangle=0$, then $\tilde{J}:Q\times\mathbb{R}\to\mathfrak{g}^{*}$ given by $\tilde{J}(q,r)=J(q)$ satisfies
@@
On the other hand, given a symplectic moment map $\tilde{J}:Q\times\mathbb{R}\to\mathfrak{g}^{*}$, we verify $\partial_{r}\langle\tilde{J},v\rangle=\imath_{\partial_{r}}\imath_{v_{Q\times\mathbb{R}}}\tilde{\omega}=p_{Q}^{*}(\imath_{v_{Q}}\eta)=0.$ Thus, $\tilde{J}$ does not depend on $r$, and we can define $J:Q\to\mathfrak{g}^{*}$ such that $\tilde{J}=p_{Q}^{*}J$. If $\tilde{\xi}$ is the unique $p_{\mathbb{R}}$-vertical vector field which is $p_{Q}$-related with $\xi$, then $\imath_{\tilde{\xi}}\imath_{v_{Q\times\mathbb{R}}}\tilde{\omega}=0.$ Therefore, $\xi\langle J,v\rangle=\tilde{\xi}\langle\tilde{J},v\rangle=0.$ Consequently, from this correspondence we conclude that the obstruction to the existence of $\tilde{J}$ is the same as $J$. The result follows from [ORbook, Prop. 4.5.17] and from the homotopy equivalence between $Q$ and $Q\times\mathbb{R}$. ∎
The proof of the previous result relies on the construction of a symplectic structure on $Q\times\mathbb{R}$, known as the symplectization procedure. As briefly mentioned, one can derive a symplectic moment map on $Q\times\mathbb{R}$ from a cosymplectic moment map on $Q$. In fact, the symplectic reduction $(Q//G\times\mathbb{R},\tilde{\omega}_{red})$ at the zero level induces the cosymplectic reduction $(Q//G,\omega_{red},\eta_{red})$ via
$\omega_{red}=i^{*}\tilde{\omega}_{red},\qquad\eta_{red}=i^{*}(\imath_{\partial r}\tilde{\omega}_{red}),$
where $i:Q//G\to Q//G\times\mathbb{R}$ is the natural inclusion $i(m)=(m,0).$
2.2. Oversymplectic groupoids
Let $\mathcal{G}\rightrightarrows M$ be a Lie groupoid endowed with a closed multiplicative 2-form $\omega\in\Omega^{2}(\mathcal{G})$, such that $\text{rank}(\omega_{x})=2\dim(M)$ for all $x\in M$. The pair $(\mathcal{G},\omega)$ is called an oversymplectic groupoid, and was originally studied in [BCWZ] in the context of integration of Dirac structures. The multiplicativity of the 2-form descends to infinitesimal data. To this end, we first recall some basic notions from the theory of infinitesimal multiplicative forms, following [BC]. Let $(A,\rho,[\cdot,\cdot])$ be a Lie algebroid, and let $\mu:A\to\Lambda^{k-1}T^{*}M$ be a vector bundle map. The map $\mu$ is called an IM $k$-form  [^1] The definition used here is actually known as closed IM forms, and it is not the most general one. However, since we will work only with closed forms, it is enough to consider this simpler case instead of the general case. if